• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 12:43:29 on February25,2009

    Entry number 263263


    keyword=Compton detector photon rates

    I checked the compton detector rates. They seem to have been at 0 since this morning. When beam came back no rates could be seen, but noticed HV current was at 1600 uA instead of the usual 1076 uA, looking on the scope no signal could be seen. I lowered the HV until I could see signal, the gain was still down by a factor 10 looking at the scope. Also looking at the scope the bacground was on the order of a few MHz in the crystal. After compensating the gain loss with an amplification of 10 and asked for lower current, I got a rate on which MCC could make the tune. And the rates were fixed by a change on prebuncher setting.

    FIGURE 1